OpenStudy (anonymous):

simplify the expression, (-6+i)/(-5+i)

OpenStudy (anonymous):

Please! I need help!!!!!! @hero

OpenStudy (anonymous):

-6 --- -5 I think thats how it goes.. Not 100%

OpenStudy (anonymous):

That isn't one of my answer choices tho, lol

OpenStudy (anonymous):

Whats the choices

OpenStudy (anonymous):

\[\frac{ -6+\iota }{ -5+\iota }\times \frac{ -5-\iota }{ -5-\iota }=\frac{ 30+6 \iota-5 \iota-\iota^2 }{ \left( -5 \right)^2-\iota^2 }=?\]

OpenStudy (anonymous):

31+1 ------ 26

OpenStudy (anonymous):

what? @surjithayer

OpenStudy (anonymous):

29+1 ----- 26

OpenStudy (anonymous):

31+i ---- 24

OpenStudy (anonymous):

31+11i ---- 26

OpenStudy (anonymous):

\[=\frac{ 30+\iota-(-1) }{ 25-(-1) }=\frac{ 31+\iota }{ 26 }\]

OpenStudy (anonymous):

so its a? @surjithayer

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks so much! @Jennifer_Thomas and @surjithayer

OpenStudy (anonymous):

yw